OpenStudy (loser66):

PDE help, please if v,w are solution of \(u_t+uu_x+u_{xxx}=0\) show that v+w is also a solution only when v*w does not depend on x

OpenStudy (loser66):

\((v+w)_t+(v+w)(v+w)_x+(v+w)_{xxx}=0\) if and only if \(vw_x+wv_x=0\) but I don't know how to turn it to vw does not depend on x

OpenStudy (loser66):

@wio

OpenStudy (kainui):

v*w does not depend on x is the same as: \[(v*w)_x = 0\]

OpenStudy (anonymous):

\[ vw_x+wv_x = (vw)_x = 0 \]

OpenStudy (anonymous):

This is the product rule

OpenStudy (loser66):

What?? that is ... amazing!! OMG.Thank you so much. I thought of something on the sky!!